when certain mass is suspended from a string,it is extended by a distance l.The string is now cut into two halves and the same mass is suspended freely from one end from one of two halves.The new extension of the spring will be,

Grade:12

1 Answers

Arun
4 years ago
 

T = 2\pi\sqrt{\frac{m}{k}}. Also Spring\; constant\; (k)\; \alpha\; \frac{1}{Length\;(l)}

When the spring is hal in length, then k becomes twice.

\therefore T` = 2\pi \sqrt{\frac{m}{2k}} \Rightarrow \frac{T`}{T} = \frac{1}{\sqrt{2}} \Rightarrow T` = \frac{T}{\sqrt{2}}

 

Time period of oscillation for spring mass system -

T= 2\pi \sqrt{\frac{m}{K}}

- wherein

m = mass of block

K = spring constant